OpenStudy (anonymous):

A racecar starting from rest accelerates uniformally at rate of 4.90 meters per second^2. What is the cars speed after it has traveled 200. meters?

OpenStudy (anonymous):

ok.., using this formula: vt^2 = vo^2 + 2 a.s; where s = 200 m, a = 4.90 m/s^2, and Vo = 0; The cars speed after is has traveled 200 is: vt^2 = 0 + (2x 4.90 x 200) vt^2 = 980 vt = sqrt (1960) vt = 44.2718 m/s Done ! :) please correct me if I wrong

OpenStudy (anonymous):

what does vo stand for?

OpenStudy (anonymous):

and vt as well?

OpenStudy (anonymous):

vo is initial velocity of racecar.., and vt is the final speed

OpenStudy (anonymous):

oh okay, thank you so much!!!

OpenStudy (anonymous):

ur welcome :)